OpenStudy (anonymous):

x^2+y-25=0 find the x and y intercepts

OpenStudy (anonymous):

x-int: y=0 solve for x y-int: x=0 solve for y

OpenStudy (anonymous):

y=25-x^2 y-intercept, plug in x=0 : y=25 x-intercept, plug in y=0 : x=5 or -5

OpenStudy (anonymous):

here is another problem y=x^2-4x-12

OpenStudy (anonymous):

do the same thing... y=(x+2)(x-6) plug in x=0: y=-12 plug in y=0: x=-2 or 6

OpenStudy (anonymous):

thankyou!

OpenStudy (anonymous):

np